280 grams of CaO is dissolved in enough water to make 10 L of solution.
1
Expert's answer
2020-11-12T06:31:47-0500

I guess your question is what the concentration of CaO in the solution

Molar mass of CaO = 40+16= 56g/mol

Mole of CaO dissolved = mass / molar mass = 280/56= 5 moles

Concentration = mole/ volume = 5mol/10L = 0.5mol/L